OpenStudy (anonymous):

How do I find the magnitude and the direction angle for vector <12√2,-12√2>?

OpenStudy (anonymous):

if i recall correctly maginitude is \[\sqrt{a^2+b^2}\] just as in pythagoras

OpenStudy (anonymous):

as for the angle, draw a picture and you will find instantly that it is \(-\frac{\pi}{4}\)

OpenStudy (anonymous):

correct so I got 16 however that is not one of the answer choices so I guessed I was doing something wrong

OpenStudy (anonymous):

\[\sqrt{(12\sqrt{2})^2+(12\sqrt{2})^2}\] i don't think it is 16

OpenStudy (anonymous):

yeah i entered in my calc wrong

OpenStudy (anonymous):

24 mag and -pi/4 is the solution then?

OpenStudy (anonymous):

\[\sqrt{288+288}=\sqrt{576}=24\]

OpenStudy (anonymous):

that is what i got, yes

OpenStudy (anonymous):

ok great thank you so much

OpenStudy (anonymous):

yw